Calculate the height through which a crane can lift a load of 4t, when its motor of 4 HP operates for 10 s. [Take,g= 10m/s^2]

Grade:10

1 Answers

Shreya D Nanda
6 years ago
Hi dear,
Before solving let us brief it out.
1HP= 746 watts ; 4HP = 2984 watts
4tonnes =4000 kg
 
coming to the question,
 
Power = work/time = force * displacement / time
according to the ques,
Power = 2984 watts
Force= mg = 4000 * 10  = 4*10^4 N
Displacement= the required height = h metres
Time = 10 s
P= F*h/t
\Rightarrow h= P* t/F
    \Rightarrow h= \left (2984 * 10 \right )/4*10^4 = 2984/4000
     \Rightarrow h= 0.746 m
    The crane can lift the load to a height of 0.746 m.
 
 Hop it is clear.....